# Enhancement of Magnon Transport by Superconductor Meissner Screening

### v1

@@ Abstract
In the calculation, we take the YIG film of thickness $1002d_{F}=$ 100 nm, saturated magnetization $\mu_{0}M_{0}=0.24$ T [35, 56], exchange stiffness $\alpha_{\rm ex}=3\times 10^{-16}~{}{\rm m}^{2}$, and damping coefficient $\alpha_{G}=10^{-4}$ [16], biased by the in-plane field $\mu_{0}H_{0}=0.05$ T. We use the superconducting NbN of thickness $d_{S}=300$ nm and London’s penetration depth $\lambda_{L}=80$ nm at $T=0.1T_{c}=1.2$ K [47, 48, 49]. Comparing the dispersions at $T=0.1T_{c}$ and $T_{c}$, Fig. 3(a) shows decreasing the temperature, the magnon dispersion is strongly shifted upwards only when $k_{y}>0$. There is no frequency shift at the ferromagnetic resonance $k_{y}=0$ due to the total reflection of stray fields by the SC, and the shift vanishes when the wave vector is large since the exchange interaction dominates. The frequency shift is then largest when $k_{y}\sim\lambda_{L}^{-1}$, resulting in the non-monotonic modulation of the group velocity as in Fig. 3(b): the group velocity is strongly enhanced (slightly suppressed) when $k_{y}\in[0,\lambda_{L}^{-1}]$ ($k_{y}\gtrsim\lambda_{L}^{-1}$). Furthermore, the lifetime of magnons is not suppressed, as shown in Fig. 3(c).
@@
According to Eq. (11), the group velocity and lifetime of magnons jointly govern their characteristic decay length. As plotted in Fig. 3(e), when decreasing the temperature from $T=T_{c}$ to $0.1T_{c}$, the decay of the excited magnetization becomes much slower due to the screening of the stray field by SCs. Figure 3(d) and (f) plot the sharp dependencies of decay length on excited frequencies $\Omega$ and temperatures, in which we find the SC generally enhances the magnon transport when the excited wave vector $k_{y}\lesssim\lambda_{L}^{-1}$. Particularly, we estimate the enhancement for YIG can be as large as $450\%$, resulting in the decay length exceeding millimeters at lower frequencies.
@@
Conclusion and discussion.—In conclusion, we develop the model calculations of the inductive interaction between magnons and SC supercurrent and predict an efficient approach to unidirectionally enhance the transport of magnons in the microwave frequencies even in the best ferrimagnet YIG via the screening of its dynamical stray field by adjacent SCs. Such mechanism is free of charge and spin-orbit-coupling, thereby also suitable for the metallic ferromagnet [57]. The frequency of some antiferromagnets [59, 58] lies in the control window as well. At room temperature, the associated mechanism relies on the perfect diamagnetism of conventional SCs for microwaves, which may be replaceable by other materials with large diamagnetic susceptibilities [60, 61]. Our finding may stimulate explorations of magnon on-chip quantum information and transport in the context of superconducting magnonics.
@@ Acknowledgements.
This work is financially supported by the National Key Research and Development Program of China under Grant No. 2023YFA1406600, the National Natural Science Foundation of China under Grants No. 12374109 and No. 12274260, as well as the startup grant of Huazhong University of Science and Technology. We thank Toeno van der Sar, Gerrit Bauer, and Yafei Ren for the useful discussions.

### v2

@@ Abstract
In the calculation, we take the YIG film of thickness $1002d_{F}=$ 100 nm, saturated magnetization $\mu_{0}M_{0}=0.24$ T [45, 61], exchange stiffness $\alpha_{\rm ex}=3\times 10^{-16}~{}{\rm m}^{2}$, and damping coefficient $\alpha_{G}=10^{-4}$ [15], biased by the in-plane field $\mu_{0}H_{0}=0.05$ T. We use the superconducting NbN of thickness $d_{S}=300$ nm and London’s penetration depth $\lambda_{L}=80$ nm at $T=0.1T_{c}=1.2$ K [48, 49, 50]. Comparing the dispersions when $T=0.1T_{c}$ and that in the absence of SCs, Fig. 4(a) shows the magnon dispersion is strongly shifted upwards only when $k_{y}>0$. There is no frequency shift at the ferromagnetic resonance $k_{y}=0$ due to the total reflection of stray fields by the SC, and the shift vanishes when the wave vector is large since the exchange interaction dominates. The frequency shift is then largest when $k_{y}\sim\lambda_{L}^{-1}$, resulting in the non-monotonic modulation of the group velocity as in Fig. 4(b): the group velocity is strongly enhanced (slightly suppressed) when $k_{y}\in[0,\lambda_{L}^{-1}]$ ($k_{y}\gtrsim\lambda_{L}^{-1}$). Furthermore, the modulation on the magnon lifetime, which is slightly (not) suppressed when $k_{y}\in[0,\lambda_{L}^{-1}]$ ($k_{y}\gtrsim\lambda_{L}^{-1}$), is much weaker than that of the group velocity, as shown in Fig. 4(c).
@@
According to Eq. (11), magnon group velocity and lifetime jointly govern their characteristic decay length. As plotted in Fig. 4(e), the decay of the excited magnetization becomes much slower due to the screening of the stray field by SCs. Figure 4(d) and (f) plot the sharp dependencies of decay length on excited frequencies $\Omega$ and temperatures, in which we find the SC generally enhances the magnon transport when the excited wave vector $k_{y}\lesssim\lambda_{L}^{-1}$. Particularly, we estimate the enhancement for YIG can be as large as $450\%$, resulting in the decay length exceeding millimeters at lower frequencies.
@@
Conclusion and discussion.—In conclusion, we develop the model calculations of the inductive interaction between magnons and SC supercurrent and predict an efficient approach to unidirectionally enhance the transport of magnons in the microwave frequencies even in the best ferrimagnet YIG via the screening of its dynamical stray field by adjacent SCs. Such a mechanism is free of charge and spin-orbit-coupling, which is also suitable for metallic ferromagnets [62] because of its relatively small conductivity. The frequency of some antiferromagnets [64, 63] also lies in the control window, and the modulation may be efficient for both two polarized modes. At room temperature, the associated mechanism relies on the perfect diamagnetism of conventional SCs for microwaves, which may be replaceable by other materials with large diamagnetic susceptibilities [65, 66]. Our finding may stimulate explorations of magnon on-chip quantum information and transport in the context of superconducting magnonics.
@@ Acknowledgements.
This work is financially supported by the National Key Research and Development Program of China under Grant No. 2023YFA1406600, the National Natural Science Foundation of China under Grants No. 12374109 and No. 12274260, as well as the startup grant of Huazhong University of Science and Technology. We thank Toeno van der Sar, Gerrit Bauer, and Yafei Ren for useful discussions.